# AGN - host galaxy photometric decomposition using a fast, accurate and precise deep learning approach

### v1

@@ 4.1 Zoobot model performance
Figure 9: Outlier fraction (calculated as the fraction of galaxies with RAE $>20\%$) as a function of the injected AGN fraction (top left panel), redshift (top right), signal-to-noise ratio (bottom left) and magnitude (bottom right). The purple lines correspond to the results from Zoobot and the green, from GALFIT. The solid lines correspond to the whole sample while the dashed and dotted lines correspond to the merger and non-merger galaxies, respectively.
@@
In Fig. 7 we show the RMSE of the Zoobot predictions (purple lines) as a function of the injected AGN fraction, redshift, signal-to-noise and magnitude (calculated in the F150W filter). The solid lines show the RMSE of the whole test subset. To separately investigate how the performance might change for highly disturbed or irregular galaxies, we also show galaxies classified as mergers (dashed lines) and galaxies classified as non-mergers (dotted lines), separately. In general, the RMSE is slightly lower for non-merging galaxies, but the difference with mergers is very small. We see that the RMSE is smaller for high and low values of the injected AGN fraction ($f_{AGN}\gtrsim 0.8$ and $f_{AGN}\lesssim 0.2$) and has a maximum mean value of 0.35 over the whole range of $f_{AGN}$. In fact at $f_{AGN}>0.6$, for mergers, and $f_{AGN}>0.4$, for non-mergers, Zoobots achieves smaller errors than the coefficient of variation of the PSF, that is, smaller error than the error on the PSF that results from the temporal and spatial variations (shaded region of Fig. 7). At intermediate levels of AGN contribution, the RMSE is a few times worse, perhaps indicating that image-based decomposition is harder when the brightness levels of the AGN and the host galaxy are comparable. There is a small increase of RMSE with increasing redshift, with a maximum value of RMSE $=0.043$ at $z=0.92$, possibly because galaxies at higher redshifts are smaller and tend to have lower signal-to-noise ratios (i.e. more dominated by noise). The latter is also observed when looking at the RMSE as a function of the signal-to-noise, where we see that the RMSE increases suddenly at $S/N<10$. Finally, there is also an increase in the RMSE with increasing magnitude (i.e. towards fainter galaxies).
@@
The purple lines (solid lines represent the whole test subset, dashed lines mergers and dotted lines non-merger) show the relative error from the Zoobot predictions as a function of the injected AGN fraction, redshift, signal-to-noise and magnitude (calculated in the F150W filter). We find that, while the RMSE decreases for low $f_{AGN}$, the relative error increases as $f_{AGN}$ decreases, particularly for $f_{AGN}<0.1$. While the RMSE only quantifies the absolute difference between real and predicted, the relative error does so in relation to the actual value. The relative error does not change with redshift. But, similarly to the RMSE, it increases with decreasing signal-to-noise and magnitude, as expected. Overall, there is a small difference between mergers (dashed lines) and non-mergers (dotted lines), with non-mergers having slightly higher relative errors in some cases.
Finally, we explore the outlier fraction in the Zoobot predictions. Here we define outliers as the Zoobot predicted AGN fractions with an RAE of more than 20%, i.e.,
@@
In Fig. 9 we show the outlier fraction as a function of the injected AGN fraction, redshift, signal-to-noise and magnitude in purple lines. The overall outlier fraction for the whole sample is 7.8%. At intermediate to high levels of AGN contribution, the outlier fraction is extremely low (close to zero). However, at low levels of injected AGN fractions, particularly for $f_{AGN}<0.1$, it increases very sharply with decreasing $f_{AGN}$. This behaviour is partly caused by our definition of outliers (i.e., it depends on whether we use a threshold on relative or absolute errors). The outlier fraction remains more or less constant as a function of redshift. Concerning the signal-to-noise ratio, there is only a small increase in the outlier fraction for $S/N<10$. At $S/N>10$, the outlier fraction drops to almost zero. As expected, there is a constant increase in the outlier fraction with decreasing magnitude. There is no significant difference in the outlier fractions between mergers (dashed lines) and non-mergers (dotted lines). In Table 4 we summarise the results obtained from Zoobot, in terms of the RMSE, RAE, mean difference and outlier fractions (at different percentage levels, 20% and 30%) for the whole test sample and for mergers and non-mergers separately.
@@ 4.2 GALFIT performance
In Fig. 10 we compare the AGN fraction obtained from running GALFIT (in the two ways explained above) with the real injected AGN fraction. When we calculate the AGN fraction by dividing by the total flux of the galaxy, we find that $\text{RMSE}=0.090$, with a mean difference (between predicted and real value) of $\langle{\Delta f_{AGN}}\rangle=-0.018$ and dispersion: $\sigma(\Delta f_{AGN})=0.086$. However, when we consider the AGN fraction calculated completely from the model, we obtain slightly worse results, with $\text{RMSE}=0.141$, $\langle{\Delta f_{AGN}}\rangle=-0.072$ and $\sigma(\Delta f_{AGN})=0.121$. This highlights how an inaccurate Sérsic model can bias the recovered AGN fraction. From now on, we only show the results from calculating the total flux of the galaxy from the original galaxy image. The RMSE obtained from GALFIT is more than three times higher than that obtained from Zoobot. There seems to be a small systematic offset (as seen by the mean difference) that is not found in the Zoobot results, and a larger spread between predicted and real (injected) AGN fractions than from Zoobot. In terms of dependence on galaxy structural properties, we find that galaxies with $n<1$, have an increased RMSE of 47%, and it is even higher when the injected AGN fraction is small ($f_{AGN}<0.2$), which means that GALFIT results are more affected by low Sérsic index than our deep learning methods. We do not see a worse prediction of the $f_{AGN}$ for galaxies with high Sérsic index for either method (see Fig. 13 in Appendix A for more details on the effect of Sérsic index in predicting $f_{AGN}$).
@@
Figure 10: Comparison between the real injected AGN fraction and the AGN fraction obtained from GALFIT fitting, obtained by dividing the PSF flux by the total flux. In the left panel, the total flux corresponds to the flux measured within a 3″aperture in the original image, while in the right panel, the flux is measured within a 3″aperture in the model (Sérsic+PSF). In both cases, the PSF flux is measured within the same aperture in the model PSF image. The dispersion in the left panel corresponds to $RMSE=0.119$, and in the right, $RMSE=0.138$. The colour bar provides the number of sources in each 2D bin.
@@
Figures 7, 8 and 9 show the RMSE, relative error and outlier fractions, respectively, as defined in the previous Section. The green lines show the results from the AGN fraction obtained from GALFIT fitting. We see similar trends to those from Zoobot, but with larger values. We see that over all the galaxy properties ($f_{AGN}$, redshift, S/N and magnitude), the RMSE is always higher for GALFIT than for Zoobot results. Similarly to the Zoobot results, we observe a small increase of RMSE with increasing redshift and with decreasing S/N. We also observe a general decrease of RMSE with increased AGN fractions, although the RMSE also decreases for $f_{AGN}<0.1$. This decrease for very low AGN fractions seems to arise due to the real AGN fraction value being very small. Actually, when looking at how the RAE changes with AGN fraction (Fig. 8) we see that it increases significantly as the AGN fraction decreases. The same results are observed with Zoobot, but again, the total RAE for GALFIT is higher than for Zoobot independently of AGN fraction, redshift, S/N or magnitude. Consequently, the outlier fraction is also always higher for GALFIT, with a dramatic increase, which was also observed in the previous Section, for $f_{AGN}<0.2$. In Table 4, we summarise the results obtained from GALFIT, for the two ways to calculate the AGN fractions, that is, by either determining the total flux of the galaxy (within 3″) from the original galaxy or the GALFIT model (Sérsic+PSF) and show the RMSE, RAE and outlier fractions (at different percentage levels, 20% and 30%) for the whole test sample.
@@
Figure 11: Number of galaxies for which GALFIT returns a good fit (blue histograms) and for which there is a bad fit or no fit from GALFIT (red histograms), as a function of the injected AGN fraction (top left panel), redshift (top right panel), signal-to-noise ratio (bottom left panel) and magnitude (bottom right panel). While there is no obvious trend with redshift, GALFIT tends to fail for bright galaxies, galaxies with low S/N, or with very high or low injected AGN fractions.
@@ 5 Summary and Conclusions
In this paper, we presented a new method to determine the AGN contribution to the total flux of a galaxy, based on DL methods and cosmological hydrodynamical simulations. We constructed the training sample for our DL model from the IllustrisTNG simulations. First, we mimicked the observational effects of JWST/COSMOS-WEB in the F150W filter, by using the same pixel scale, convolving with survey PSFs, adding Poisson noise and finally, adding real JWST background. Then, we artificially added an AGN component (approximated by the PSF) at different levels to represent different AGN contributions to the total flux. After training our model we analysed the results in detail on a test sample. Furthermore, we used GALFIT to fit the light profile of the galaxies in the test sample with a combination of Sérsic + PSF profiles to calculate the AGN contribution within a circular aperture. We showed how our DL method outperforms the traditional 2D surface brightness fitting method. Our main results are summarised in the following:
@@
    The $f_{AGN}$ predicted by Zoobot has a very tight correlation with the injected $f_{AGN}$, with a mean difference of $-0.0006$ and an overall RMSE $=0.027$ for the whole test sample, demonstrating how well our DL method is at recovering the true AGN contribution to the total flux. This performance metric is three times better than that obtained using GALFIT. In both methods, there is a decrease in the RMSE at low and high $f_{AGN}$. However, the low RMSE at $f_{AGN}\lesssim 0.2$ is, in part, a natural consequence of very small values of the true/injected $f_{AGN}$. In both methods, we see an increase in RMSE as redshift increases. While for Zoobot results the RMSE increases towards low S/N and high magnitudes, that is not the case for GALFIT, for which the RMSE remains roughly constant as a function of S/N and magnitude.
@@
    The outlier fraction given by Zoobot, as measured by the number of galaxies that have RAE $>20\%$, is below 8% for most of the sample, except for galaxies with very low AGN fractions ($f_{AGN}<0.1$) and faint galaxies (mag $>22$). There is only a small increase with magnitude and lower S/N. However, the outlier fraction is most affected by the AGN fractions and starts to increase for $f_{AGN}<0.2$. The outlier fraction that we obtain from the GALFIT results is on average around 20%, and similarly to our Zoobot results, there is an increase with decreasing $f_{AGN}$, S/N and increasing magnitude. Furthermore, there is a small increase in the outlier fraction at higher redshift, which is not observed in the Zoobot results.
@@
    In constructing the training sample for our DL model, we fully incorporated the temporal and spatial variations of the PSF, which helps our model to better determine the intrinsic AGN fraction. And once it is trained our DL model can be easily applied to new data, without the need for user inputs as 2D fitting codes (such as GALFIT) do. Furthermore, it can output $f_{AGN}$ predictions of thousands of galaxies in a few seconds, making it an ideal method for large extra-galactic imaging surveys.
In future works, we will study how AGNs selected via other methods, such as via X-ray, mid-infrared, or radio observations, compare with AGNs identified using our DL method. We can investigate in detail what is the $f_{AGN}$ distribution of the X-ray/mid-infrared/radio-selected AGNs. A follow-up and complementary approach that we also aim to develop in the near future will focus on not only inferring the level of AGN contribution to the observed total light but also decomposing the observed images into a pure AGN component and a host galaxy component. This will allow us to be able to investigate host galaxies of AGNs in more detail. For example, we can look for the presence of merging signs and bar features which could trigger AGNs.

### v2

@@ 4.1 Zoobot model performance
Figure 9: Outlier fraction (calculated as the fraction of galaxies with RAE $>20\%$) as a function of the injected AGN contribution fraction (top left panel), redshift (top right), S/N (bottom left) and $r$ (bottom right). The purple lines correspond to the results from Zoobot and the green, from GALFIT. The solid lines correspond to the whole sample while the dashed and dotted lines correspond to the merger and non-merger galaxies, respectively.
@@
In Fig. 7 we show the RMSE of the Zoobot predictions (purple lines) as a function of the injected AGN contribution fraction, redshift, S/N and kron radius ($r_{kron}$, calculated with SExtractor in the F150W filter). The solid lines show the RMSE of the whole test subset. To separately investigate how the performance might change for highly disturbed or irregular galaxies, we also show galaxies classified as mergers (dashed lines) and galaxies classified as non-mergers (dotted lines), separately. In general, the RMSE is slightly lower for non-merging galaxies, but the difference with mergers is very small. The RMSE is larger for lower values of the injected AGN contribution fraction and reaches a maximum mean value of 0.020. At $f_{AGN}>0.2$, Zoobot achieves smaller errors than the coefficient of variation of the PSF, that is, smaller error than the error on the PSF resulting from the intrinsic temporal and spatial variations (shaded region of Fig. 7). In other words, the precision of our DL-based method is better than the intrinsic variations of the observed PSF. This is only possible because the DL method is trained on data which include the full range of observed PSFs. There is an increase in the RMSE with increasing redshift, with a maximum value of RMSE $=0.026$ at $z=3$, possibly because galaxies at higher redshifts are smaller and tend to have lower S/N (i.e. more dominated by noise). Indeed, we see that generally the RMSE increases with decreasing $r_{kron}$ and S/N.
@@
The purple lines (solid lines represent the whole test subset, dashed lines mergers and dotted lines non-merger) show the relative error from the Zoobot predictions as a function of the injected AGN contribution fraction, redshift, S/N and $r_{kron}$ (calculated by SExtractor in the F150W filter). We find that the relative error increases as $f_{AGN}$ decreases, particularly for $f_{AGN}<0.1$. While the RMSE only quantifies the absolute difference between real and predicted, the relative error does so in relation to the actual value. The relative error does not change much with redshift. But, similarly to the RMSE, it increases with decreasing S/N and decreasing $r_{kron}$, as expected. Overall, there is a small difference between mergers (dashed lines) and non-mergers (dotted lines), with non-mergers having slightly higher relative errors in some cases.
Finally, we explore the outlier fraction in the Zoobot predictions. Here we define outliers as the Zoobot predicted AGN contribution fractions with an RAE of more than 20%, that is,
@@
In Fig. 9, we show the outlier fraction as a function of the injected AGN contribution fraction, redshift, S/N and $r_{kron}$ in purple lines. The overall outlier fraction for the whole sample is 6.5%. At intermediate to high levels of AGN contribution, the outlier fraction is extremely low (close to zero). However, at low levels of injected AGN fractions, particularly for $f_{AGN}<0.1$, it increases very sharply with decreasing $f_{AGN}$. This behaviour is partly caused by our definition of outliers (i.e., it depends on whether we use a threshold on relative or absolute errors). The outlier fraction remains more or less constant as a function of redshift and galaxy size. Concerning the S/N, there is only a small increase in the outlier fraction for $S/N<3$. At $S/N>3$, the outlier fraction drops to almost zero. There is no significant difference in the outlier fractions between mergers (dashed lines) and non-mergers (dotted lines).
In Table 3 we summarise the results obtained from Zoobot, in terms of the RMSE, RAE, mean difference and outlier fractions (at different percentage levels, 20% and 30%) for the whole test sample and for mergers and non-mergers separately.
@@ 4.2 GALFIT performance
In Fig. 10 we compare the AGN contribution fraction obtained from running GALFIT (in the two ways explained above) with the real injected AGN contribution fraction. When we calculate the AGN contribution fraction by dividing by the total flux of the galaxy, we find that $\text{RMSE}=0.12$, with a mean difference (between predicted and real value) of $\langle{\Delta f_{AGN}}\rangle=-0.018$ and dispersion: $\sigma(\Delta f_{AGN})=0.12$. However, when we consider the AGN contribution fraction calculated completely from the model, we obtain slightly worse results, with $\text{RMSE}=0.17$, $\langle{\Delta f_{AGN}}\rangle=-0.075$ and $\sigma(\Delta f_{AGN})=0.15$. This highlights how an inaccurate Sérsic model can bias the recovered AGN contribution fraction. From now on, we only show the results from calculating the total flux of the galaxy from the original galaxy image. The RMSE obtained from GALFIT is more than 10 times higher than that obtained from Zoobot. There seems to be a small systematic offset (as seen by the mean difference) that is not found in the Zoobot results, and a larger spread between predicted and real (injected) AGN contribution fractions than from Zoobot. In terms of dependence on galaxy structural properties, we find that galaxies with $n<1$, have an increased RMSE of 34%, and it is even higher when the injected AGN contribution fraction is small ($f_{AGN}<0.2$), which means that GALFIT results are more affected by low Sérsic index than our deep learning methods. We do not see a worse prediction of the $f_{AGN}$ for galaxies with high Sérsic index for either method (see Fig. 16 in Appendix A for more details on the effect of Sérsic index in predicting $f_{AGN}$).
@@
Figure 10: Comparison between the real injected AGN contribution fraction and the AGN contribution fraction obtained from GALFIT fitting, obtained by dividing the PSF flux by the total flux. In the left panel, the total flux corresponds to the flux measured within a 2″aperture in the original image, while in the right panel, the flux is measured within a 2″aperture in the model (Sérsic+PSF). In both cases, the PSF flux is measured within the same aperture in the model PSF image. The dispersion in the left panel corresponds to $RMSE=0.12$, and in the right, $RMSE=0.17$. The colour bar provides the number of sources in each 2D bin.
@@
Figure 7, Fig. 8 and Fig. 9 show the RMSE, relative error and outlier fractions, respectively, as defined in the previous section. The green lines show the results from the AGN contributino fraction obtained from the GALFIT fitting. We see similar trends to those from Zoobot, but with larger values (indicating significantly worse performance). We see that over all the galaxy properties ($f_{AGN}$, redshift, S/N and $r_{kron}$) considered, the RMSE is always higher for GALFIT than for Zoobot. Similar to the Zoobot results, we observe a small increase of RMSE with increasing redshift and with decreasing S/N. We also observe a general decrease in the RMSE with increased AGN contribution fractions, although the RMSE also decreases at $f_{AGN}<0.1$. This decrease for very low AGN contribution fractions seems to arise due to the real AGN fraction value being very small. Actually, when looking at how the RAE changes with AGN contribution fraction (Fig. 8) we see that it increases significantly as $f_{AGN}$ decreases. The same results are observed with Zoobot, but again, the total RAE for GALFIT is higher than for Zoobot, independently of $f_{AGN}$, redshift, S/N or $r_{kron}$. Consequently, the outlier fraction is also always higher for GALFIT, with a dramatic increase, which was also observed in the previous section, at $f_{AGN}<0.2$. In Table 3, we summarise the results on the performance metrics obtained from GALFIT, for the two ways to calculate the AGN contribution fractions, that is, by either determining the total flux of the galaxy (within 2″) from the original galaxy or the GALFIT model (Sérsic+PSF) and show the RMSE, RAE and outlier fractions (at different percentage levels, 20% and 30%) for the whole test sample.
We find that $24\pm 1\%$ of all the galaxies from the subsample test set have no or bad fit from GALFIT. However, there is not a significant difference in the fraction of mergers and non-mergers that have bad or no fit. In Fig. 11, we show that GALFIT fails more often for galaxies with high or low AGN contribution fraction (particularly at $f_{AGN}\gtrsim 0.8$ and $f_{AGN}\lesssim 0.2$), higher redshift, lower S/N and very small galaxies. Therefore, even though we have previously seen that both the RMSE and RAE are relatively low at $f_{AGN}\gtrsim 0.8$, GALFIT fails to find a good fit in that regime for $46\pm 3\%$ of the galaxies, while at $f_{AGN}\lesssim 0.2$ GALFIT fails for $25\pm 2\%$ of galaxies. GALFIT also fails more often at the higher redshifts ($28\pm 1\%$ of fails for $z>1$), compared to $19\pm 1\%$ for the rest of the sample, and $14\pm 2\%$ in the lowest redshift bin. Finally, GALFIT is more likely to fail for faint galaxies (with low S/N) and for very bright ones (with high S/N). There are $75\pm 3\%$ of galaxies with no good fit for $S/N>8$ and $24\pm 4\%$ for $S/N<3$. These results highlight the importance of good-quality imaging in order to find a good fit to the galaxy’s light with GALFIT, while our DL-based method can always determine a galaxy’s AGN contribution fraction with good accuracy and precision regardless of these galaxy properties (within the dynamic ranges tested in this work).
Figure 11: Number of galaxies for which GALFIT returns a good fit (blue histograms) and for which there is a bad fit or no fit from GALFIT (red histograms), as a function of the injected AGN contribution fraction (top left panel), redshift (top right panel), $S/N$ (bottom left panel) and $r_{kron}$ (bottom right panel). GALFIT tends to fail for bright galaxies, galaxies with low S/N, or with very high or low injected AGN contribution fractions.
4.3 Application to JWST galaxies
@@
Finally, we apply the trained DL model to our stellar mass complete sample of real JWST galaxies as described in Section 2.1. In Fig. 12, we show the distribution of the DL predicted AGN contribution fractions. Considering the overal RMSE of our DL-based method in recovering the intrinsic $f_{AGN}$ is 0.013, a $5\sigma$ cut corresponding to $f_{AGN}=0.065$ can be used to select galaxies for which the method is confident in identifying a non-negligible contribution from a central point source component. Indeed, the distribution shown in Fig.12 clearly separates at around this $5\sigma$ cut value. Note that many traditional methods of selecting AGN use a binary approach to separate galaxies into those which host AGN and those which do not. In reality, there is a continuous distribution of AGN contribution fraction. Therefore, although a binary AGN vs non-AGN selection approach is useful in identifying galaxies with dominant AGN, it is limited in the sense that it does not reflect the full spectrum. Adopting the $5\sigma$ cut at $f_{AGN}=0.065$, we find a total of $3337\begin{subarray}{c}+218\\ -196\end{subarray}$ galaxies hosting a non-negligile level of AGN activity in the 0.28 deg2 of the COSMOS field analysed in this paper. Therefore, with respect to the total size of our stellar mass complete sample, 13% of the galaxies are identified by our method to have a non-neglible level of AGN contribution in the JWST/NIRCAM F150W filter.
Using the continuous parameter $f_{AGN}$ predicted by our DL method, users can choose different cuts on AGN contribution fraction to construct an AGN sample depending on the specific requirements. For example, we can classify galaxies as AGN based on the fractional AGN contribution, using a cut at $f_{AGN}>0.2$, which leads to $937\begin{subarray}{c}+111\\ -170\end{subarray}$ galaxies being classified as AGN over 0.28 deg2. With a less conservative cut at $f_{AGN}>0.1$, we find $2611\begin{subarray}{c}+358\\ -753\end{subarray}$ AGN within our sample. Furthermore, we find that there are 18 galaxies with $f_{AGN}>0.7$ (shown in Fig. 17 in the appendix). There are five galaxies with spectroscopic information within those 18 galaxies. Four out of those five (i.e. 80%) show optical broad lines in their spectra, as indicated by the specific flag in the release catalogue of spectroscopic redshifts available to the COSMOS collaboration from KECK follow up observations (Khostovan et al 2025, in prep). Based on this, we conclude that most galaxies with high $f_{AGN}$, as predicted by our method, are consistent with optical spectroscopically confirmed AGN (when that information is available)
Next, we compare our AGN based on the estimated $f_{\rm AGN}$ with AGN samples selected via X-ray detection or MIR colours. Using the cut at $f_{\rm AGN}>0.2$, we find that 20 % of the X-ray AGN are also selected as AGN based on the estimated $f_{\rm AGN}$, while 8 % of the MIR selected AGN are also selected by our method. If we use a less conservative cut at $f_{\rm AGN}>0.1$, then the overlapping fractions increase for both AGN selections, to 15 % overlap with the MIR-selected AGN and 33 % overlap with the X-ray AGN. There is, however, a large fraction of the X-ray and MIR AGN with very low AGN contribution fraction $f_{\rm AGN}<0.1$, and therefore, are not selected as AGN by our DL method based on a 0.1 threshold on $f_{AGN}$. This is to be expected, as it is well-known that some X-ray and MIR AGN are not detected by optical diagnostics. This phenomena could be explained by a bright host galaxy that dilutes the light of the AGN or by the presence of copious dust in the host galaxy, particularly when the galaxy is viewed edge-on, causing heavy obscuration (Satyapal et al., 2008; Jackson et al., 2012; Smith et al., 2014; Hickox & Alexander, 2018; Fitriana & Murayama, 2022). In Fig. 18 and 19 we show a random subset of the X-ray and MIR AGN with $f_{\rm AGN}<0.065$ (i.e., below the $5\sigma$ cut based on the overall RMSE of our DL method), respectively. Many of these galaxies display spiral, clumpy or edge-on morphologies, in which higher dust content is expected. And in all cases, they do not show any obvious nuclear component in the JWST/NIRCAM F150W images, which explains the low predicted $f_{\rm AGN}$ despite being classified as X-ray or MIR AGN.
In Fig. 13 we show the normalised distributions of the predicted $f_{\rm AGN}$ in the different AGN samples. The X-ray-selected AGN tend to dominated at higher AGN contribution fractions $f_{\rm AGN}>0.1$, indicating better correspondence with optically dominant AGN. For comparison, we also plot the distribution of the predicted $f_{\rm AGN}$ in the ‘non-AGN’ sample which correspond to galaxies not identified as X-ray or MIR AGN in our sample. We can see that while a large fraction (around 90 %) of the non-AGN have $f_{\rm AGN}<0.1$, a small fraction (around 4 %) of them do have predicted $f_{\rm AGN}>0.2$, indicating that they can be AGN missed by the X-ray and MIR selections. In fact, it is widely known that the X-ray-selected AGN can miss a significant fraction of AGN with strong optical emission lines, particularly those that are heavily absorbed (Heckman et al., 2005), while the MIR selection can miss low-luminosity AGN or AGN in host galaxies dominated by starburst, or radio-loud AGN (Hainline et al., 2016; Truebenbach & Darling, 2017). Out of the 18 galaxies with $f_{AGN}>0.7$ from our DL-based method as discussed previously, only a third of them are identified as MIR or X-ray AGN. However, we expect most of those 18 galaxies would be optically dominant AGN, based on the fact that five out of the 18 galaxies were followed up in optical spectroscopy and four exhibit broad lines in their spectra.
The top panel of Fig. 14 shows the overlapping fraction of AGN also identified using our DL method based on the predicted AGN contribution fraction as a function of the adopted cut on the X-ray luminosity $L_{X}$ or the rest-frame 6 $\mu m$ luminosity $L_{6\mu m}$, for the X-ray and MIR selected AGN, respectively. We see that the overlapping fraction of AGN also selected by our model increases with increasing luminosity (more slowly for the MIR AGN). This is expected as the X-ray luminosity and the rest-frame 6 $\mu m$ luminosity (which are indicators of the power of the X-ray or MIR-selected AGN) correlate, albeit with significant scatter, with the luminosity of the AGN component in the JWST/NIRCAM F150W filter $L_{\rm AGN}$ (calculated as the total luminosity in that filter multiplied by $f_{\rm AGN}$). We show these correlations in the bottom panel of Fig. 14, only for galaxies with $f_{\rm AGN}>0.065$ for which we can confidently identify a central point source contribution). The strongest correlation is found for the X-ray AGN detected in the soft band, and the weakest correlation is found in the MIR AGN.
Figure 13: Fraction of galaxies in bins of predicted AGN contribution fraction in the X-ray AGN, MIR AGN and ’non-AGN’ samples.
Figure 14: Top panel: Overlapping fraction of AGN also identified by our DL method as a function of the adopted cut on the X-ray or the rest-frame 6 $\mu$m luminosity for the X-ray and MIR-selected AGN. The solid lines show the overlapping fraction by applying a cut at $f_{\rm AGN}>0.2$, while the dashed lines show the overlapping fraction if we select AGN by requiring $f_{\rm AGN}>0.1$. Bottom panel: AGN luminosity in the JWST/NIRCAM F150W filter versus the X-ray luminosity or the rest-frame 6 $\mu m$ luminosity, for the X-ray-selected or MIR-selected AGN. The dash-dotted lines show the respective linear fits.
@@ 5 Summary and Conclusions
In this paper, we presented a new DL-based method Zoobot to determine the AGN contribution fraction to the total flux of a galaxy, trained on realistic mock images generated from cosmological hydrodynamical simulations. Specifically, we constructed the training sample for our DL model from the IllustrisTNG simulations. First, we mimicked the observational effects of theJWST/COSMOS-WEB survey in the NIRCAM F150W filter, by using the same pixel scale, convolving with the full range of the survey PSFs, adding Poisson noise and finally, adding the real JWST sky background. Then, we artificially added an AGN component (approximated by the PSF) at different levels to represent different AGN contribution fractions to the total flux. After training our model we analysed the performance results in detail on a test sample. Furthermore, we used GALFIT to fit the 2D light profile of the galaxies in the test sample with a combination of Sérsic + PSF profiles to calculate the AGN contribution fraction within a circular aperture. We showed how our DL method significantly outperforms the traditional 2D surface brightness fitting method, in precision, accuracy, speed and success rate. Our main results are summarised in the following:
@@
    The AGN contribution fraction $f_{AGN}$ predicted by Zoobot has a very tight correlation with the injected $f_{AGN}$, with a mean difference of $-0.0018$ and an overall RMSE $=0.013$ for the whole test sample over the redshift range $30.5<z<$ 3, demonstrating how well our DL method is at recovering the true AGN contribution to the total flux. The RMSE performance metric is around 10 times better than that obtained using GALFIT. In both methods, there is an increase in the RMSE at lower $f_{AGN}$ values, higher redshifts, lower S/N and smaller galaxy sizes.
@@
    The outlier fraction given by Zoobot, as measured by the number of galaxies that have RAE $>20\%$, is below 8% for most of the sample, except for galaxies with very low AGN contribution fractions ($f_{AGN}<0.1$) and faint galaxies (mag $>22$ in the F150W filter). There is only a small increase with lower S/N. However, the outlier fraction is most affected by the AGN contribution fractions and starts to increase sharply at $f_{AGN}<0.2$. The outlier fraction obtained from the GALFIT results is on average around 20%, and similarly to our Zoobot results, there is an increase with decreasing $f_{AGN}$ and S/N. Furthermore, there is a significant increase in the outlier fraction at higher redshift, which is not observed in the Zoobot results.
@@
    In constructing the training sample for our DL model, we fully incorporated the temporal and spatial variations of the PSF, which helps our model to better determine the intrinsic AGN contribution fraction. At $f_{AGN}>0.2$, the precision of our DL method is even better than the level of intrinsic variations in the observed PSF. Once it is trained, our DL model can be easily applied to new data, without the need for user inputs as in traditional 2D fitting codes (such as GALFIT) do. Furthermore, it can output $f_{AGN}$ predictions of thousands of galaxies in a few seconds, making it an ideal method for large extra-galactic imaging surveys.
  4. iv)
    We applied the trained DL model to real JWST observations, and we find that $937\begin{subarray}{c}+111\\ -170\end{subarray}$ have $f_{AGN}>0.2$, while $2611\begin{subarray}{c}+358\\ -753\end{subarray}$ have $f_{AGN}>0.1$. Moreover, when comparing with other AGN selection methods, we find that 20% of the X-ray AGN and 8% of the MIR selected AGN are also selected as AGN based on the selection of $f_{\rm AGN}>0.2$. Using the selection of $f_{\rm AGN}>0.1$, the overlapping fractions increase to 33% and 15% for X-ray and MIR AGN, respectively.
In future works, we will extend the comparison of our deep learning predictions with other AGN selection techniques, such as those identified using optical spectroscopy (type 1 and type 2 optical AGN) and radio observations. A follow-up and complementary approach that we also aim to develop in the near future will focus on not only inferring the level of AGN contribution fraction to the observed total light but also decomposing the observed images into a pure AGN component and a host galaxy component. This will allow us to investigate host galaxies of AGN in more detail. For example, we can look for the presence of merging signs and bar features which could trigger AGN.